OpenStudy (anonymous):

Help! The sample space of an experiment is {A, B, C}, where P(A) + P(B) = 0.71, P(B) + P(C) = 0.44. Find P(A), P(B), and P(C).

OpenStudy (anonymous):

P(A)+P(B)+P(C)+P(B)=0.71+0.44=1.15 1+P(B)=1.15 P(B)=1.15-1=0.15 Now you can find P(A) and P(C)

OpenStudy (anonymous):

Where did you get the 1 from?

OpenStudy (anonymous):

total probability=1 \[P(A)+P(B)+P(C)=1\]

OpenStudy (anonymous):

Got it =) Thank You